# MUDDFormer: Breaking Residual Bottlenecks in Transformers via Multiway Dynamic Dense Connections

### v1

@@ C.2 Baseline Models
DenseFormer: The DenseFormer model from Pagliardini et al. (2024) without dilation, which has the best performance according to the paper. We implemented the model in JAX based on the PyTorch code released by the authors666https://github.com/epfml/DenseFormer.
@@
Appendix D Image Classification with ViT
Besides decoder-only transformer for language modeling, we apply MUDD connections to Vision Transformer (ViT, an encoder-only Transformer) (Dosovitskiy et al., 2020) for image classification on the ImageNet-1k dataset (ILSVRC-2012). Implementation and experimental settings (e.g. the use of RandAugmention+MixUp, fixed 2D sincos position embedding and global average pooling) are based on the Big Vision codebase777https://github.com/google-research/big_vision. We use AdamW with = 0.9, = 0.999, gradient clip value of 1.0, weight decay of 0.3, learning rate of 0.003 with 10000 warmup steps followed by cosine decay to 0, batch size of 4096, RandAugment of 1evel 10, Mixup of 0.2 and dropout rate of 0.1. We use ViT-S/16 as the baseline model and equip it with MUDD connections to obtain MUDDViT-S/16. We also compare with a 1.72× larger model ViT-M/16 (Table 6). We report validation loss and top-1 accuracy results on 90 and 300 epochs in Table 7. As can be seen, the gain from MUDD connections decreases a bit during the training progress, probably because many epochs of repeated passes over the same dataset diminish the additional expressive capacity brought by MUDD connections. Despite this, MUDDViT-S/16 still outperforms ViT-S/16 by 2% on epoch 300, also surpassing ViT-M/16.
@@ Appendix D Image Classification with ViT
Appendix E Visualization
Head activation from attention patterns In Section 3.3, we show that the ratio of head activations in MUDDPythia-2.8B is larger than that in Pythia-2.8B. Here we draw the actual attention patterns on a randomly sampled sequence of length 32 from Pile validation set for the 32 heads in layer 25 of these two models in Figure 9 and 10. It is clear that attentions in Pythia mainly concentrate on the sink token (inactive) while attentions in MUDDPythia disperse on various tokens (active).
Cross-layer dynamic weights To better understand MUDDPythia, we visualize dynamic dense connection weights. Due to the high variance of the norm of hidden states , we scale those weights by the average norm of each layer, rectifying the importance of weights. The rectified mean and standard deviation of dynamic connection weights in MUDDPythia-2.8B are shown in Figure 11 and 12, respectively. It is evident that the patterns of the four streams (query, key, value, residual) differ from each other, validating the necessity of separating them from the standard residual stream. It is noteworthy that in the value-stream connections, most layers have a salient and more dynamic weight on output of the first layer, thus forming a long-range channel to transport bottom information for attention heads in upper layers.

### v2

@@ D.2 Baseline Models
DenseFormer: The DenseFormer model from Pagliardini et al. (2024) without dilation, which has the best performance according to the paper. We implemented the model in JAX based on the PyTorch code released by the authors888https://github.com/epfml/DenseFormer.
@@
Appendix E Memory Usage and Wall-Clock Time for Main Experiments
Theoretical analysis of memory usage In theory, peak activation memory usage for training a transformer in float16 with layers, heads, sequence length , batch size and model dim occurs at the beginning of backpropagation and is composed of two parts:
1.
hidden states for layers: (gradient checkpointing)
2.
activation memory for a layer: (outlined in (Korthikanti et al., 2023), recomputation of layer when backpropagate it)
We compare theoretical activation memory usages in Table 6. DenseFormer adds to store gradients for each layer’s hidden state when backpropagating DA after layer . Based on this, MUDDFormer adds another for recomputation of layer ’s multi-way hidden states ,, when backpropagate it (they are not stored for all layers but are recomputed during backpropagation). The extra memory ratio for MUDD is , and typical values are less than 30%. During inference, the activation memory usage is dominated by the cache, which is not impacted by the extra memory brought by MUDD.
model | activation memory | memory ratio |
TFM++ | 1 | |
DenseFormer | ||
MUDDFM |
Measured memory usage and wall-clock time In Table 7, we report actual memory usage (measured using jax.profile) and wall-clock time for both the main (Figure 3, Table 3) and efficiency (Table 4) experiments. Besides activation memory, the actual memory usage also includes model parameters, gradients and optimizer states, and is affected by JAX compiler optimizations. For all model architectures and sizes, the relative training speed is 80%-90%, which could be further improved by custom kernel implementation. The extra memory ratio of MUDD is 20%-30%, comparable to that of HyperConnections (see Table 9 in their paper). As noted in our paper, the results for efficiency experiments (row 3-5) are of more practical relevance because they represent a more commonly used architecture (Transformer++) and model sizes.
model | model size | wall-clock time (hour) | rel. speed | mem (GB) | mem ratio | v5p pod size | tokens | batch size |
TFM++ / MUDDFM | 405M | 5.7/7 | 81% | 86/111 | 29% | 16 | 7B | 0.5M |
834M | 20/25.1 | 79% | 225/273 | 21% | 16 | 15B | 0.5M | |
1.4B | 29.5/32.5 | 91% | 301/386 | 28% | 32 | 26B | 0.5M | |
2.8B | 122/145 | 84% | 1352/1648 | 22% | 128 | 300B | 2M | |
6.9B | 251/262 | 96% | 1887/2222 | 17% | 128 | 300B | 2M | |
Pythia / MUDDPythia | 1.4B | 163/183 | 89% | 1296/1655 | 28% | 64 | 300B | 2M |
2.8B | 124/154 | 81% | 1318/1655 | 25% | 128 | 300B | 2M |
Appendix F Image Classification with ViT
Besides decoder-only transformer for language modeling, we apply MUDD connections to Vision Transformer (ViT, an encoder-only Transformer) (Dosovitskiy et al., 2020) for image classification on the ImageNet-1k dataset (ILSVRC-2012). Implementation and experimental settings (e.g. the use of RandAugmention+MixUp, fixed 2D sincos position embedding and global average pooling) are based on the Big Vision codebase999https://github.com/google-research/big_vision. We use AdamW with = 0.9, = 0.999, gradient clip value of 1.0, weight decay of 0.3, learning rate of 0.003 with 10000 warmup steps followed by cosine decay to 0, batch size of 4096, RandAugment of 1evel 10, Mixup of 0.2 and dropout rate of 0.1. We use ViT-S/16 as the baseline model and equip it with MUDD connections to obtain MUDDViT-S/16. We also compare with a 1.72× larger model ViT-M/16 (Table 8). We report validation loss and top-1 accuracy results on 90 and 300 epochs in Table 9. As can be seen, the gain from MUDD connections decreases a bit during the training progress, probably because many epochs of repeated passes over the same dataset diminish the additional expressive capacity brought by MUDD connections. Despite this, MUDDViT-S/16 still outperforms ViT-S/16 by 2% on epoch 300, also surpassing ViT-M/16.
@@ Appendix F Image Classification with ViT
Appendix G Visualization
Head activation from attention patterns In Section 3.4, we show that the ratio of head activations in MUDDPythia-2.8B is larger than that in Pythia-2.8B. Here we draw the actual attention patterns on a randomly sampled sequence of length 32 from Pile validation set for the 32 heads in layer 25 of these two models in Figure 10 and 11. It is clear that attentions in Pythia mainly concentrate on the sink token (inactive) while attentions in MUDDPythia disperse on various tokens (active).
Cross-layer dynamic weights To better understand MUDDPythia, we visualize dynamic dense connection weights. Due to the high variance of the norm of hidden states , we scale those weights by the average norm of each layer, rectifying the importance of weights. The rectified mean and standard deviation of dynamic connection weights in MUDDPythia-2.8B are shown in Figure 12 and 13, respectively. It is evident that the patterns of the four streams (query, key, value, residual) differ from each other, validating the necessity of separating them from the standard residual stream. It is noteworthy that in the value-stream connections, most layers have a salient and more dynamic weight on output of the first layer, thus forming a long-range channel to transport bottom information for attention heads in upper layers.